The sum of a number,and 5 less than double the number is 34. What is the number

Answer by Timnewman(323)

Let the number be "x"
The sum of a number and 5 less than double the number reads:
2x-5+x
if the result is 34,
then we write,
2x-5+x=34
3x-5=34
3x=34+5
3x=39
x=13
The number is 13.
Hope this helps;..
----TIMNEWMN------